OpenStudy (swagster23780):

Ben wants to join a fitness club. The fitness club charges an initial membership fee of $49.50 and a monthly fee of $17.50. Part A Ben has $172 to spend on a membership at the fitness club. How many months can Ben be a member of the fitness club? Ben has enough money to be a member of the fitness club for months. Part B Ben’s company will pay $300.00 each year toward his fitness club membership. Ben thinks he will have to pay $40.50 for his membership his first year. Is he correct? Use the drop-down menus to explain your reasoning.

rebeccaxhawaii (rebeccaxhawaii):

17.50x+49.50=172

rebeccaxhawaii (rebeccaxhawaii):

17.50x=172-49.50 17.50x = 122.50

rebeccaxhawaii (rebeccaxhawaii):